# DeSTA2.5-Audio: Toward General-Purpose Large Audio Language Model with Self-Generated Cross-Modal Alignment

### v1

@@ Abstract
We introduce DeSTA2.5-Audio, a general-purpose Large Audio Language Model (LALM) designed for robust auditory perception and instruction-following, without requiring task-specific audio instruction-tuning. Recent LALMs typically augment Large Language Models (LLMs) with auditory capabilities by training on large-scale, manually curated or LLM-synthesized audio-instruction datasets. However, these approaches have often suffered from the catastrophic forgetting of the LLM’s original language abilities. To address this, we revisit the data construction pipeline and propose DeSTA, a self-generated cross-modal alignment strategy in which the backbone LLM generates its own training targets. This approach preserves the LLM’s native language proficiency while establishing effective audio-text alignment, thereby enabling zero-shot generalization without task-specific tuning. Using DeSTA, we construct DeSTA-AQA5M, a large-scale, task-agnostic dataset containing 5 million training samples derived from 7,000 hours of audio spanning 50 diverse datasets, including speech, environmental sounds, and music. DeSTA2.5-Audio achieves state-of-the-art or competitive performance across a wide range of audio-language benchmarks, including Dynamic-SUPERB, MMAU, SAKURA, Speech-IFEval, and VoiceBench. Comprehensive comparative studies demonstrate that our self-generated strategy outperforms widely adopted data construction and training strategies in both auditory perception and instruction-following capabilities. Our findings underscore the importance of carefully designed data construction in LALM development and offer practical insights for building robust, general-purpose LALMs.  [^1] https://github.com/kehanlu/DeSTA2.5-Audio
Index Terms:
Large Audio Language Model, cross-modal alignment, dataset construction, instruction-tuning
@@ I Introduction
To develop a general-purpose LALM, two core capabilities are essential: auditory perception and instruction-following. Auditory perception refers to the comprehensive processing of auditory information, including speech, non-verbal cues, background sounds, and music. In addition, instruction-following entails interpreting diverse user commands, integrating them with internal knowledge, and producing contextually appropriate responses. To achieve this, most LALM architectures adopt a modular approach, combining a pre-trained audio model with a text-based LLM as the knowledge backbone [8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. These components are integrated through a cross-modal alignment process, which typically involves joint fine-tuning on a corpus of paired audio and text instructions to bridge the representational gap between the audio and text modalities. Accordingly, the research problem centers on training the fusion model with limited audio-related data, while leveraging the LLM’s generalization capabilities to handle unseen tasks.
@@
However, acquiring high-quality data for cross-modal alignment remains a significant challenge. A common strategy involves transforming existing audio datasets by converting them into audio instruction-tuning datasets, consisting of (1) audio input, (2) textual instructions, and (3) textual responses. These training triplets can be generated either by text-based LLMs or through manual annotation. To improve the generalization capabilities of LALMs, several techniques have been adopted, including fine-tuning lightweight LoRA adapters[11, 10, 12, 14, 13], adopting multi-stage curriculum learning [12, 16, 17, 13, 14, 15, 20], and applying test-time pruning [12]. While current LALMs perform well across a range of audio-related benchmarks [28, 29, 30, 31, 32, 33, 34], many studies highlight the issue of catastrophic forgetting [8, 9, 12, 11, 10, 14, 35, 36], where the models become overly specialized to a limited set of audio-related tasks at the cost of generalization to unseen tasks. Moreover, our previous work, Speech-IFEval [36], also reveals that most LALMs experience significant degradation in text-based knowledge, underscoring the persistent limitations of current cross-modal alignment techniques.
To address the issue of catastrophic forgetting in cross-modal alignment, we revisit the data construction process with a key observation: annotations from different sources, such as various LLMs or human annotators, exhibit distinct stylistic and behavioral patterns. While this may seem intuitive, prior studies often rely on responses generated by powerful LLMs [11, 10, 16, 17, 13], such as GPT [1], or human-designed pairs [12, 14] to ensure high-quality supervision. However, we hypothesize that incorporating annotations from heterogeneous sources inevitably introduces biases into the dataset. As a result, the fusion model must not only learn auditory grounding but also implicitly adapt to the new response patterns introduced by the annotators. We believe this extra adaptation burden can gradually distort the LLM’s native output distribution or instruction-following behavior, ultimately degrading its text-based knowledge.
@@
Building on this insight, we introduce DeSTA [^2] DeSTA stands for Descriptive Speech-Text Alignment, first introduced in [8] and further improved with self-generated design in [9]. While the method generalizes beyond speech in this study, we retain the original naming convention for consistency., a self-generated cross-modal alignment framework that mitigates conflicting supervision by having the backbone language model generate its own training targets. Specifically, we convert the metadata of each audio segment into a structured textual description and pair it with an arbitrary prompt. Then, the LLM generates a corresponding response, which serves as the training target for cross-modal alignment. This self-generated supervision ensures stylistic and semantic consistency with the LLM’s native output distribution, thereby preserving its instruction-following capabilities while enabling effective adaptation to auditory inputs. In our prior work, DeSTA2 [9], we demonstrate that training solely on a single general task, speech captioning, is sufficient for learning robust audio grounding. The resulting model exhibited strong zero-shot generalization across a wide range of downstream tasks without any task-specific instruction tuning, presenting a scalable path toward general-purpose LALM training. In this work, we extend DeSTA by scaling our dataset construction pipeline across diverse audio domains, including speech, environmental sounds, and music. We aggregate 50 publicly available audio datasets spanning a wide range of audio information to create a large-scale corpus of approximately 5 million audio-text pairs, which we call DeSTA-AQA5M. This dataset encompasses various forms of audio information, including paralinguistic cues, speaker identities, audio quality indicators, and environmental or contextual sounds. This expanded model, referred to as DeSTA2.5-Audio, further validates the effectiveness of self-generated alignment in enabling general and robust audio-language modeling.
We evaluate DeSTA2.5-Audio on several audio-language benchmarks, including Dynamic-SUPERB [33, 32], MMAU [28], SAKURA [29], Speech-IFEval [36], and VoiceBench [30]. Despite using only 7,000 hours of training data, compared to 510,000 hours used by the baseline Qwen2-Audio-Instruct [17], our model achieves state-of-the-art or competitive results across domains. Our comparative analysis also emphasizes the critical role of the distribution mismatch problem. Models trained on self-generated targets consistently outperform those trained on mismatched data sources (e.g., from other LLM), even when the latter originate from larger or more powerful LLMs. Although adaptation techniques like LoRA [37] can alleviate distribution mismatches, they often come at the cost of reduced performance in complex reasoning or out-of-domain scenarios. These results suggest that training data distribution should be a fundamental design consideration in the development of LALMs.
@@
  • We propose DeSTA, a self-generated cross-modal alignment method that leverages the backbone LLM to generate its own training targets, eliminating the need for task-specific instruction-tuning. By reducing the mismatch between training data and the model’s original behavior, DeSTA prevents the catastrophic forgetting problem.
@@
  • Our model, DeSTA2.5-Audio, achieved superior results across multiple audio-language benchmarks. We demonstrate strong performance on both auditory perception and instruction-following tasks, showing significantly enhanced robustness compared to existing LALMs.
  • We conducted extensive comparison studies validating that self-generated data is critical for success, revealing significant limitations in existing data construction approaches. Our analysis demonstrates that using different LLMs for data generation leads to suboptimal performance or model collapse, and that even stronger models from the same family do not guarantee improvements, providing important insights and design considerations for future research.
@@ II-A Large Audio Language Model (LALM)
There are already several overview papers about LALMs; Therefore, we briefly review only the technologies directly relevant to this work. For a comprehensive overview, please refer to these overview papers [38, 39, 40, 41, 34].
@@ II-B Self-Generation Mechanism
The concept of self-generated data has gained significant traction in text LLM fine-tuning [50, 51, 52]. Rather than relying on supervision from a powerful teacher model or human-annotated data, self-generated approaches enable the model to create its own training data. A key advantage of this method is that the training targets are naturally aligned with the model’s inherent behavior and data distribution. Several studies on text LLM have shown that this alignment can improve performance [53, 54, 55, 56, 57].
In multi-modal contexts, researchers have begun to adapt this approach to the speech domain. One common method uses automatic speech recognition (ASR) transcriptions as inputs and allows the backbone LLM to generate textual continuations as training targets [58, 59, 60, 61]. However, these efforts are often limited to spoken-content-centric tasks, such as ASR and spoken question answering, and tend to underutilize the rich, multidimensional information present in audio. Moreover, these studies rarely analyze how model behavior or performance might differ if the training data were not self-generated.
Our prior work, DeSTA2 [9], extended the self-generated paradigm in the speech domain by incorporating paralinguistic cues, speaker characteristics, and other non-verbal audio signals. We found that training on self-generated “speech captions,” which provide comprehensive audio grounding, enables the model to yield strong zero-shot performance on a variety of tasks, demonstrating the ability to generalize from a single training objective. In this work, we broaden our approach to encompass a wider variety of audio information, and we provide comprehensive evaluations and analyses to validate its effectiveness.
@@ III Methodology
An overview of the DeSTA framework is shown in Figure 2. Just as humans can imagine auditory scenes from written narratives, such as voices and ambient sounds in a novel or film script, modern LLMs exhibit a similar abilities. Most auditory concepts, such as tone, speaking style, and environmental sounds, can be described in natural language. Based on this intuition, DeSTA aims to provide an approximate textual description of audio input, which will be used to generate the corresponding training target using the backbone LLM.
@@ III-A Self-Generated Dataset Construction
During the data construction phase, we first collect a diverse set of datasets containing detailed metadata. The metadata associated with each audio segment is converted into a structured textual format following the schema below:
@@
For example, a speech clip can be represented as “[00:00-00:05] Hello world (Gender:Female, Emotion:Happy...)”, while an audio captioning sample could appear as “[00:00-00:10] (A dog barking)”. This conversion allows us to represent diverse audio types using a consistent textual format. The resulting initial dataset, denoted as $\mathcal{D}_{\text{initial}}=\{({x}^{\text{audio}},{x}^{\text{text}})\}$, aligns each audio clip ${x}^{\text{audio}}$ with its corresponding textual description ${x}^{\text{text}}$.
Based on these initial audio–description pairs, we utilize a text-based LLM to generate training targets that reflect how the model would naturally respond to the given descriptions. Specifically, we randomly sample a text prompt $\mathbf{p}$ from a predefined instruction pool $\mathcal{P}$, which includes diverse prompt types such as descriptive tasks (e.g., “Describe the audio”), role-playing scenarios (e.g., “Respond to the audio based on its expression”), and open-ended questions (e.g., “Where is the audio being recorded?”). These prompts are crafted to maximize the LLM’s ability to utilize all available information from the textual description, thereby enhancing the quality of cross-modal alignment. The backbone LLM then takes the text description $x^{\text{text}}$ and the prompt $\mathbf{p}$ as inputs and produces a response $\mathbf{y}=\text{LLM}(x^{\text{text}},\mathbf{p})$. This automated setup enables the generation of detailed and context-aware responses, eliminating the need for labor-intensive, task-specific instruction design. Through this procedure, we construct an audio instruction-tuning dataset in the form of $\mathcal{D}={(x^{\text{audio}},x^{\text{text}},\mathbf{p},\mathbf{y})}$, where each data point includes the audio input, its corresponding text description, the sampled prompt, and the LLM-generated response.
A key advantage of this method is that it preserves the LLM’s native interpretation and response behavior to the inputs, ensuring stylistic and semantic consistency in the training data. For example, we observe that the instruction-tuned Llama3.1 [4] typically generates explanatory responses, organizes content with bullet points, and often includes greeting phrases before the main content. These stylistic patterns, which are specific to the LLM, are naturally embedded in the generated training data. Thus, while the construction pipeline is compatible with any text-based LLM, employing the same model, i.e., the self-generated setting, is the most reasonable design for cross-modal alignment.
@@ III-B Model Training
The audio input $x^{\text{audio}}$ is first encoded into a continuous representation by applying Q-Former blocks [24] to multiple intermediate hidden states from the audio encoder. Formally, let $\mathbf{h}^{(\ell)}\in\mathbb{R}^{T\times d}$ denote the output from the $\ell$-th encoder layer, where $T$ represents the temporal dimension and $d$ is the hidden dimension. For each selected layer $\ell$, we define learnable query vectors $\mathbf{Q}^{(\ell)}\in\mathbb{R}^{N\times d}$, where $N$ is the number of queries. The Q-Former computes:
@@
The outputs from multiple layers are aggregated using learnable scalar weights $\alpha^{(\ell)}$ with the constraint $\sum_{\ell}\alpha^{(\ell)}=1$, and subsequently projected through a linear layer to match the LLM embedding dimension $d^{\prime}$:

### v2

@@ Abstract
We introduce DeSTA2.5-Audio, a general-purpose Large Audio Language Model (LALM) designed for robust auditory perception and instruction-following. Recent LALMs augment Large Language Models (LLMs) with auditory capabilities by training on large-scale audio-instruction datasets. However, existing LALMs have often suffered from the catastrophic forgetting of the LLM’s original abilities. Therefore, balancing knowledge retention and audio perception has become a critical challenge. To address this, we revisit the data construction pipeline and propose a self-generated cross-modal alignment strategy in which the backbone LLM generates its own training targets, named DeSTA. This approach aims at preserving the LLM’s native language proficiency thereby enabling zero-shot generalization without task-specific tuning. We construct DeSTA-AQA5M, a large-scale, task-agnostic dataset containing 5 million training samples derived from 7,000 hours of audio spanning 50 diverse datasets, including speech, environmental sounds, and music. DeSTA2.5-Audio achieves state-of-the-art or competitive performance across a wide range of audio-language benchmarks, including Dynamic-SUPERB, MMAU, SAKURA, Speech-IFEval, and VoiceBench. Comprehensive comparative studies demonstrate that our self-generated strategy outperforms existing training strategies. Our findings underscore the importance of carefully designed data construction in LALM development and offer practical insights for building robust, general-purpose LALMs.  [^1] https://github.com/kehanlu/DeSTA2.5-Audio
@@ I Introduction
To develop a general-purpose LALM, two core capabilities are essential: auditory perception and instruction-following. Auditory perception refers to the comprehensive processing of auditory information, including speech, non-verbal cues, background sounds, and music. In addition, instruction-following entails interpreting diverse user commands, integrating them with internal knowledge, and producing contextually appropriate responses. To achieve this, most LALMs adopt a modular architecture, combining a pre-trained audio model with a text-based LLM as the knowledge backbone [72, 73, 35, 36, 92, 34, 47, 61, 18, 17, 1, 50, 33, 42]. These components are integrated through a cross-modal alignment process, which involves joint fine-tuning on a corpus of paired audio and text instructions to bridge the representational gap between the audio and text modalities. Accordingly, the research problem centers on training the fusion model with limited audio-related data, while leveraging the LLM’s generalization capabilities to handle unseen tasks.
@@
However, acquiring high-quality data for cross-modal alignment remains a significant challenge. A common strategy involves transforming existing audio datasets by converting them into audio instruction-tuning datasets, consisting of (1) audio input, (2) textual instructions, and (3) textual responses. These training triplets can be generated either by text-based LLMs or through manual annotation. While current LALMs perform well across a range of audio-related benchmarks [86, 117, 15, 102, 49, 48, 118], many studies highlight the issue of catastrophic forgetting [72, 73, 92, 36, 35, 47, 38, 74], where the models become overly specialized to a limited set of audio-related tasks at the cost of generalization to unseen tasks. Moreover, our previous work, Speech-IFEval [74], also reveals that most LALMs experience significant degradation in text-based knowledge. Therefore, balancing text-based knowledge retention and audio generalization becomes a key challenge.
To mitigate catastrophic forgetting in LALM development, we revisit the paradigm of data construction process. We observe that a primary cause of degradation may stem from heterogeneous training targets. Specifically, previous studies often rely on training pairs generated by powerful LLMs [36, 84, 35, 18, 17, 34, 13] or human-designed pairs [92, 47]. While straightforward, we observe that these approaches inevitably introduce a distribution shift regarding the backbone LLM’s native behavior, such as distinct stylistic and behavioral patterns. Consequently, training LALM on such heterogeneous data compels the model to simultaneously learn auditory perception and stylistic patterns from the foreign annotator. This implicit adaptation interferes with the cross-modal alignment objective, diverting the model’s focus and gradually overwriting its inherent text knowledge.
@@
Building on this insight, we introduce DeSTA [^2] DeSTA stands for Descriptive Speech-Text Alignment, first introduced in [72] and further improved with self-generated design in [73]. While the method generalizes beyond speech in this study, we retain the original naming convention for consistency., a self-generated cross-modal alignment framework that mitigates conflicting supervision by having the backbone language model generate its own training targets. Specifically, we convert the metadata of each audio segment into a structured textual description and pair it with an arbitrary prompt. Then, the LLM generates a corresponding response, which serves as the training target for cross-modal alignment. This self-generated supervision ensures stylistic and semantic consistency with the LLM’s native output distribution, thereby preserving its instruction-following capabilities. In our prior work, DeSTA2 [73], we demonstrate that training solely on a single general task, speech captioning, is sufficient for learning robust audio grounding. The resulting model exhibited strong zero-shot generalization across a wide range of downstream tasks without any task-specific instruction tuning. In this work, we extend DeSTA by scaling our dataset construction pipeline across diverse audio domains, including speech, environmental sounds, and music. We aggregate 50 publicly available audio datasets spanning a wide range of audio information to create a large-scale corpus of approximately 5 million audio-text pairs, which we call DeSTA-AQA5M. This dataset encompasses various forms of audio information, including paralinguistic cues, speaker identities, audio quality indicators, and environmental or contextual sounds. This expanded model, referred to as DeSTA2.5-Audio, further validates the effectiveness of self-generated alignment in enabling general and robust audio-language modeling.
We evaluate DeSTA2.5-Audio on several audio-language benchmarks, including Dynamic-SUPERB [48, 49], MMAU [86], SAKURA [117], Speech-IFEval [74], and VoiceBench [15]. We also conduct CLAP-style analysis to study the learned “modality-gap” at the representation-level [25, 112]. Despite using only 7,000 hours of training data, compared to 510,000 hours used by the baseline Qwen2-Audio-Instruct [17], our model achieves state-of-the-art or competitive results across domains. Our comparative analysis also emphasizes the critical role of the distribution mismatch problem. Models trained on self-generated targets consistently outperform those trained on mismatched data sources (e.g., from other LLM), even when the latter originate from larger or more powerful LLMs. Although adaptation techniques like LoRA [45] can alleviate distribution mismatches, they often come at the cost of reduced performance in complex reasoning or out-of-domain scenarios. These results suggest that training data distribution should be a fundamental design consideration in the development of LALMs.
@@
  • We propose DeSTA, a self-generated cross-modal alignment method that reduces the mismatch between training data and the model’s original behavior. We prevent the catastrophic forgetting problem and eliminate the need for task-specific instruction-tuning.
@@
  • Our model, DeSTA2.5-Audio, achieved superior results across multiple audio-language benchmarks. We demonstrate strong performance on both auditory perception and knowledge-intensive tasks, showing significantly enhanced robustness compared to existing LALMs.
  • We conducted extensive comparison studies validating that self-generated design is critical for success, revealing significant limitations in previous LALMs. Our findings reshape the design philosophy and provide important insights for future research.
@@ II-A Large Audio Language Model (LALM)
There are already several overview papers about LALMs; Therefore, we briefly review only the technologies directly relevant to this work. For a comprehensive overview, please refer to these survey papers [8, 53, 81, 20, 118].
@@ II-B Self-Generation Mechanism
The concept of self-generated data has gained significant traction in text LLM fine-tuning [105, 43, 68]. Rather than relying on supervision from a powerful teacher model or human-annotated data, self-generated approaches enable the model to create its own training data. A key advantage of this method is that the training targets are naturally matched with the model’s inherent behavior and data distribution. Several studies on text LLM have shown that this alignment can improve performance [85, 67, 122, 16, 6].
In multi-modal contexts, researchers have begun to adapt this approach to the speech domain. One common method uses automatic speech recognition (ASR) transcriptions as inputs and allows the backbone LLM to generate textual continuations as training targets [27, 103, 124, 44]. However, these efforts are often limited to spoken-content-centric tasks, such as ASR and spoken question answering, and tend to underutilize the rich, multidimensional information present in audio. Our prior work, DeSTA2 [73], extended the self-generated paradigm in the speech domain by incorporating paralinguistic cues, speaker characteristics, and other non-verbal audio signals. We found that training on self-generated “speech captions,” which provide comprehensive audio grounding, enables the model to yield strong zero-shot performance on a variety of tasks, demonstrating the ability to generalize from a single training objective. Moreover, these studies rarely analyze how model behavior or performance might differ if the training data were not self-generated. In this work, we further broaden our approach to encompass a wider variety of audio information, and we provide comprehensive evaluations and analyses to validate its effectiveness.
In summary, while using text metadata to synthesize audio-text training pairs has been widely adopted in existing LALM studies, our approach differs fundamentally in design philosophy. Compared to existing LALMs such as LTU [36], SALMONN [92] and Qwen-Audio [18, 17] that rely on instruction-tuning datasets generated by heterogeneous sources, our design prioritizes a more streamlined and robust alignment strategy. We not only simplify the training pipeline, but also fundamentally avoid the trade-off between auditory perception and textual knowledge retention.
TABLE I: Data examples from the dataset construction stage. We feed the audio description and prompt to generate training targets. These examples also show that identical inputs lead to different model responses across LLMs, highlighting the distribution mismatch problem we explored in this paper.
Audio Description: [00:00-00:03] If not me, then who?
(Emotion: Sad, Gender: Female, Pitch: Normal, Volume: Very low, Speaking speed: Slow, Duration: 3s)
Prompt: Describe the core of the clip with brevity.
Model Response (Training Target)
$\bullet$ Llama3.1: A melancholic, introspective tone, questioning one’s identity and purpose.
$\bullet$ Qwen2.5: A sad female voice asks, ”If not me, then who?” in a very low, slow, and normal-pitched tone.
$\bullet$ Gemma3: A woman expresses a feeling of sadness and questioning her worth.
Audio Description: [00:00-00:10] (A helicopter vibrates loudly.)
Prompt: Describe this audio’s character.
Model Response (Training Target):
$\bullet$ Llama3.1: The audio clip is a loud, vibrating sound, likely caused by the rotor blades of a helicopter in operation.
$\bullet$ Qwen2.5: The audio features the character of a loud, vibrating helicopter.
$\bullet$ Gemma3: Loud, vibrating, and sounds like a helicopter.
@@ III Methodology
An overview of the DeSTA framework is shown in Figure 2. Just as humans can imagine auditory scenes from written narratives, such as voices and ambient sounds in a novel or film script, modern LLMs exhibit similar abilities. Most auditory concepts, such as tone, speaking style, and environmental sounds, can be described in natural language. Based on this intuition, DeSTA aims to provide an approximate textual description of audio input, which will be used to generate the corresponding training target using the backbone LLM.
@@ III-A Self-Generated Dataset Construction
During the data construction phase, we first collect a diverse set of datasets containing detailed metadata. As illustrated in Table I, we empirically convert the metadata into a structured textual format following the schema below:
@@
The resulting initial dataset, denoted as $\mathcal{D}_{\text{initial}}=\{({x}^{\text{audio}},{x}^{\text{text}})\}$, pairs each audio clip ${x}^{\text{audio}}$ with its corresponding textual description ${x}^{\text{text}}$.
Based on these initial audio–description pairs, we utilize the backbone LLM to generate training targets that reflect how the model would naturally respond to the given descriptions. Specifically, we randomly sample a text prompt $\mathbf{p}$ from a predefined instruction pool $\mathcal{P}$, which includes diverse prompt types such as descriptive tasks (e.g., “Describe the audio”), role-playing scenarios (e.g., “Respond to the audio based on its expression”), and open-ended questions (e.g., “Where is the audio being recorded?”). These prompts are crafted to maximize the LLM’s ability to utilize all available information from the textual description, thereby enhancing the quality of cross-modal alignment. The backbone LLM then takes the text description $x^{\text{text}}$ and the prompt $\mathbf{p}$ as inputs and produces a response $\mathbf{y}=\text{LLM}(x^{\text{text}},\mathbf{p})$. This automated setup enables the generation of detailed and context-aware responses, eliminating the need for labor-intensive, task-specific instruction design. Through this procedure, we construct an audio instruction-tuning dataset in the form of $\mathcal{D}={(x^{\text{audio}},x^{\text{text}},\mathbf{p},\mathbf{y})}$, where each data point includes the audio input, its corresponding text description, the sampled prompt, and the LLM-generated response.
@@ III-B Model Training
The audio input $x^{\text{audio}}$ is first encoded into a continuous representation by applying Q-Former blocks [66] to multiple intermediate hidden states from the audio encoder [72, 73]. Formally, let $\mathbf{h}^{(\ell)}\in\mathbb{R}^{T\times d}$ denotes the output from the $\ell$-th encoder layer, where $T$ represents the temporal dimension and $d$ is the hidden dimension. For each selected layer $\ell$, we define learnable query vectors $\mathbf{Q}^{(\ell)}\in\mathbb{R}^{N\times d}$, where $N$ is the number of queries. The Q-Former computes:
@@
The outputs from multiple layers are aggregated using learnable scalar weights $\alpha^{(\ell)}$ with the constraint $\sum_{\ell}\alpha^{(\ell)}=1$ [120], and subsequently projected through a linear layer to match the LLM embedding dimension $d^{\prime}$:
@@
The key advantage of the self-generated design in DeSTA is that it directs the optimization process to focus exclusively on cross-modal alignment. As an illustration, the optimization process for the trainable parameters can be conceptualized in two objectives:
  • Cross-Modal Grounding: maps audio features to the LLM’s semantic space
  • Stylistic Adaptation: shifts the output distribution to match the target’s specific phrasing and style
When training targets are derived from external sources (e.g., human annotators or different LLMs), they inevitably diverge from the backbone LLM’s native distribution, introducing a divergent training signal that gradually distorts the model behavior. In contrast, by employing the backbone LLM itself to generate training targets, we eliminate this stylistic mismatch. This interpretation aligns with our empirical findings and provides a rationale for the robust generalization and instruction-following abilities observed in DeSTA2.5-Audio.